# HD2-SSC: High-Dimension High-Density Semantic Scene Completion for Autonomous Driving

### v1

@@ Appendix A Appendix
Table 5 below presents the efficiency evaluation results of our HD2-SSC approach against the best comparison method SGN (Mei et al. 2024). It can be observed that HD2-SSC achieves superior performance with slightly increased model parameters and GFLOPs, even less GPU memory and inference time, since our enhanced feature enables accurate SSC prediction with (128,128,16) feature grids, avoiding SGN’s costly up-sampling to (256,256,32) grids.
@@
Table 6 below presents the evaluation results on the Occ3D-nuScenes (Tian et al. 2024) dataset, where our HD2-SSC outperforms popular OccFormer (Zhang, Zhu, and Du 2023) and BEVDet4D (Huang and Huang 2022) baselines, demonstrating our generalizability across different datasets.
@@
Table 7 presents the evaluation results with EfficientNet-B7 (Tan and Le 2019) backbone on the SSCBench-KITTI-360 (Li et al. 2023d) test set, where our HD2-SSC achieves superior performance compared to CGForemr (Yu et al. 2024) and L2COcc (Wang et al. 2025), further demonstrating our effectiveness in bridging the dimension and density gaps in the SSC task.

### v2

@@ Appendix A Appendix
Table 5 below presents the efficiency evaluation results of our HD2-SSC approach against the best comparison method SGN (mei2024camera). It can be observed that HD2-SSC achieves superior performance with slightly increased model parameters and GFLOPs, even less GPU memory and inference time, since our enhanced feature enables accurate SSC prediction with (128,128,16) feature grids, avoiding SGN’s costly up-sampling to (256,256,32) grids.
@@
Table 6 below presents the evaluation results on the Occ3D-nuScenes (tian2024occ3d) dataset, where our HD2-SSC outperforms popular OccFormer (zhang2023occformer) and BEVDet4D (huang2022bevdet4d) baselines, demonstrating our generalizability across different datasets.
@@
Table 7 presents the evaluation results with EfficientNet-B7 (tan2019efficientnet) backbone on the SSCBench-KITTI-360 (li2023sscbench) test set, where our HD2-SSC achieves superior performance compared to CGForemr (yu2024context) and L2COcc (wang2025l2cocc), further demonstrating our effectiveness in bridging the dimension and density gaps in the SSC task.